OpenStudy (anonymous):

The area of a rectangle is the product of width and length. If a rectangle has length 12m√21m and width √16m which of the following is the area of the rectangle?

OpenStudy (anonymous):

a=l*b \[a=12\sqrt{21} *\sqrt{16}\]

OpenStudy (anonymous):

can you give me option

OpenStudy (anonymous):

I dont have options

OpenStudy (anonymous):

\[A=12\sqrt{21}\times \sqrt{16}\] \[=12\sqrt{21}\times4\] \[=48\sqrt{21}\] metres Does that make sense?